Determine the energies in eV of the fourth and fifth energy levels of the hydrogen atom.

(a) fourth energy level
eV

(b) fifth energy level
eV
Answer #1

Here,

as the energy of energy level of Hydrogen atom is given as

E = -13.6/n^2 eV

a) for fourth energy level

n = 4

E = -13.6/4^2 eV

E = -0.85 eV

the energy of fourth level is -0.85 eV

b)

for the fith energy level

n = 5

E = -13.6/5^2 eV

E = -0.544 eV

the energy of fourth level is -0.544 eV